# Sigma: Scalable Spectral Insights for LLM Collapse

### v1

@@ Appendix A Mathematical Deductions and Proof of Theorems
$\lVert\hat{C}_{t}-C\rVert_{op}=O_{p}(t^{-1/2})$ (A.12)
@@
$\lVert\Delta_{t}\rVert_{op}\leq\lVert\hat{C}_{t}^{-1}\rVert_{op}\lVert\hat{C}_{t}-C\rVert_{op}\lVert C^{-1}\rVert_{op}.$ (A.14)
@@
This appendix contains the derivations that connect the theoretical bounds in Section 3.2 to the computable diagnostics in Algorithm 1.
@@ B.1 Regularized $\log$-determinant and PSD Monotonicity
since all eigenvalues are strictly positive after the $\delta I_{m}$ shift.
B.2 Diagnostic I: from $\beta_{k}$ to a Computable Budget $\widehat{\beta}_{k}$
Since $G_{B}^{(k)}\succeq 0$,
$\beta_{k}:=\lambda_{\max}(G_{B}^{(k)})\leq\mathrm{tr}(G_{B}^{(k)}).$ (B.2)
@@ B.2 Diagnostic I: from $\beta_{k}$ to a Computable Budget $\widehat{\beta}_{k}$
The above energy cap and tail budget gives us the PSD Majorant (Deterministic Envelope). Indeed, from $G^{(k)}=G_{A}^{(k)}+G_{B}^{(k)}$ and Ky-Fan/Weyl dominance,
@@
Shifting by $\delta I_{m}$ and applying PSD monotonicity yields the valid envelope gives us
$\log\det(G^{(k)}+\delta I_{m})\leq\log\det\!\left(G_{A}^{(k)}+(\widehat{\beta}_{k}+\delta)I_{m}\right)=:\mathcal{U}_{\mathrm{KF}}^{(k)}(\delta).$ (B.6)
Now let $\mathcal{U}^{(k)}_{\mathrm{KF},\star}(\delta):=\log\det(G_{A}^{(k)}+(\beta_{k}+\delta)I_{m})$. Because $\widehat{\beta}_{k}\geq\beta_{k}$, we obtain
@@
Rearrange the above inequality and we arrive at the conceptual normalization inequality, which tells us the the gap is controlled by the ideal (incomputable) envelope.
@@
Motivated by the deduction above, we can define the normalized reporting form as
$\mathcal{G}_{\mathrm{KF}}^{(k)}(\delta)$ $:=\mathcal{U}_{\mathrm{KF}}^{(k)}(\delta)-m\log(\widehat{\beta}_{k}+\delta),$ (B.9)
which matches (3.16).
B.3 Diagnostic II: LLN Scaling, Cancellation of $K$, and Covariance Normalization
From the LLN scaling route, we can set
@@ B.3 Diagnostic II: LLN Scaling, Cancellation of $K$, and Covariance Normalization
If $\log\det(G^{(k)})\leq\log K+\mathcal{U}_{\mathrm{LLN}}^{(k)}(0)$ for an unknown finite $K$, then baseline differencing removes $\log K$:
@@
This means that we effectively remove the unknown constant $K$ from the diagnostic. Finally, we can define the covariance-normalized form
$\mathcal{U}_{\mathrm{LLN,cov}}^{(k)}(\delta):=\mathcal{U}_{\mathrm{LLN}}^{(k)}(\delta)-m\log n_{k}=\log\det(G_{A}^{(k)}+\delta I_{m})-m\log n_{A},$ (B.12)
which is (3.17) and is the statistic computed in Algorithm 1.
@@ B.4 Numerical Computation of $\log$-determinant
Appendix C Supplementary Experiments, Reproducibility, and Extended Diagnostics (FIN)
C.1 Section Roadmap and Artifact Integrity
This section of the appendix serves three purposes.
(i) Reproducibility.
We provide a complete specification of the experimental pipeline (training, synthetic regeneration, evaluation, and Sigma-UB computation) in a set of ten self-contained tables (Appendix C.2). The intent is that an external group can rerun the entire loop without relying on implicit conventions.
(ii) Plot–table consistency.
[…]
(iii) Additional diagnostics and interpretation.
Because Sigma-UB is a geometric diagnostic (not a surface-form metric), we also report lightweight surface-form proxies—distinct-2 and hashed $n$-gram HHI—computed on the same model outputs used for Sigma-UB. We include a compact overlay plot that normalizes all signals to $[0,1]$ to visually assess co-movement (Appendix C.4). We then add a bucket-localization slice (optional) to show where contraction is most pronounced under the frozen prompt taxonomy (Appendix C.5). Throughout, we emphasize that surface-form proxies are contextual and not collapse definitions.
Note on filenames. For continuity with earlier drafts, some exported artifact filenames retain the prefix tech_ (e.g., tech_s1_two_tracks.pdf). In this appendix and in the main text, these files correspond to the FIN run described in Section 4.
C.2 Complete reproducibility specification for S1 and S2
@@
We summarize every operational choice that materially affects outcomes: model hyperparameters; synthetic regeneration and mixture; decoding; seeds; Sigma-UB computation; evaluation metrics; hardware/software; file layout; and the formal difference between S1 and S2. Unless explicitly stated, S1 and S2 share the same configuration; the only intended difference is whether weights are reset to $M^{(0)}$ (S1) or carried forward $M^{(g-1)}$ (S2).
@@
Table 2: Table 1: Model and training hyperparameters. These hyperparameters are held fixed across S1 and S2 so that observed differences are attributable to data recursion vs. data+weight recursion, rather than to optimization budget changes.
@@
Generations reported $g=0,\dots,50$ (51 checkpoints) Main text FIN run shown in Figures 12
@@
Table 3: Table 2: Data generation and mixing. We use constant-budget mixtures and, for the reported runs, set $\alpha=0$ to isolate synthetic recursion dynamics without real-data refresh.
@@
Max tokens (evaluation) 256 Prompt-bank responses Used to obtain sufficiently long responses for embedding/geometry
Table 4: Table 3: Text generation (decoding) parameters. Decoding is frozen across checkpoints to prevent evaluation drift and to keep the generation distribution comparable over time.
Aspect Specification
Base seed 42
Domain offsets Domain-specific offsets are derived via a deterministic hash (to avoid accidental seed reuse across domains).
Per-context determinism Synthetic generations are seeded deterministically per real-context ID and generation index.
Per-sample determinism When multiple sampling steps exist (e.g., batching), the seed schedule ensures reproducible ordering and reproducible stochastic draws.
Implementation note Seed utility functions are included in the codebase; reruns with the same config reproduce identical synthetic pools and identical evaluation outputs.
Table 5: Table 4: Seeds and determinism. The recursion loop is extremely sensitive to small sampling changes; therefore we enforce determinism at the level of domain, generation, and context ID.
@@
Table 6: Table 5: Sigma-UB computation. Both tracks are computed from the same frozen embeddings and the same observed index set, so drift is attributable to model evolution rather than to evaluation sampling.
@@
Excluded metrics No perplexity We intentionally do not report PPL in this revision
Table 7: Table 6: Evaluation metrics. The appendix focuses on geometry diagnostics and lightweight surface-form proxies; perplexity-based baselines are excluded by design.
@@
Table 8: Table 7: Computational environment. Hardware does not change the definition of metrics but can affect throughput; determinism is enforced via fixed seeds and fixed decoding.
Artifact Path template / convention
S1 metrics CSV <METRICS_ROOT>/final/<DOMAIN>/metrics_paper*.csv
S2 metrics CSV <METRICS_ROOT>/true_recursion/<DOMAIN>/ metrics_paper_recursion*.csv
Bucket metrics CSV (optional) <METRICS_ROOT>/final/<DOMAIN>/ metrics_buckets_paper*.csv
Main-text figures figures/<prefix>_s1_two_tracks.pdf, figures/<prefix>_s2_two_tracks.pdf
Point tables tables/<prefix>_s1_two_tracks_points.tex (and S2 analog)
Summary table tables/tab_<prefix>_s1s2_summary.tex
Surface proxy plot figures/<prefix>_surface_proxies.pdf
Table 9: Table 8: Data paths and file structure. This table documents the exact expected locations of input metrics and the produced figures/tables used by the paper.
Setting Defining update rule
S1 (restart-from-base) $M^{(g,d)}\leftarrow\mathrm{Finetune}\!\big(M^{(0)},\,D^{(g,d)}(\alpha)\big)$. Weights do not compound across generations; only the synthetic data distribution evolves.
S2 (true recursion) $M^{(g,d)}\leftarrow\mathrm{Finetune}\!\big(M^{(g-1,d)},\,D^{(g,d)}(\alpha)\big)$. Both synthetic data and weights compound across generations.
Table 10: Table 9: Key difference between S1 and S2. This is the core experimental intervention: the regeneration protocol and mixture rule are held fixed while weight carryover is toggled.
@@
Table 11: Table 10: Prompt buckets for localization. Bucket IDs and assignments are fixed by DecayBench; slicing by bucket does not change the evaluation distribution over time, it only reveals where contraction is concentrated within a frozen taxonomy.
C.3 Exact per-generation values for the main-text drift plots
Figures 12 are generated from exported per-generation tables. We include them here to (a) make the plots auditable and (b) guarantee that all reported summary scalars (final drift and OLS slope) can be recomputed from a single canonical source. In particular, the S1 vs. S2 summary in Table 1 is computed directly from these per-generation series (including the baseline point $g=0$, where drift is 0 by definition).
@@ C.3 Exact per-generation values for the main-text drift plots
Interpreting Track-I/Track-II disagreement.
A recurring pattern in S2 is that Track II indicates strong contraction while Track I does not “collapse” in the same way. This is not a contradiction: Track II measures observed-spectrum log-volume contraction (trend-sensitive), while Track I is an upper-envelope controlled by a worst-case tail-energy budget. When the tail-energy bound is loose relative to the observed sub-Gram spectrum, Track I can be non-alarming even when Track II contracts sharply. In practice, we treat the separation itself as informative: it flags “early collapse” regimes where the observed spectrum has contracted but the worst-case envelope remains permissive.
C.4 Surface-form proxy diagnostics and alignment with Sigma-UB
Sigma-UB is explicitly designed to track representation geometry rather than surface repetition. Nevertheless, it is useful to ask whether geometry contraction coincides with familiar surface-form failure modes. We therefore compute two lightweight proxies on the same generated text used for Sigma-UB embeddings: (i) distinct-2 (unique/total bigrams; higher is more diverse), and (ii) hashed $n$-gram HHI (higher is more concentrated/repetitive). These proxies are not collapse definitions and can be noisy, but they provide an operational sanity check.
Figure 3: Normalized overlay of Sigma-UB drifts and surface-form proxies (FIN). Each curve is min–max normalized to $[0,1]$ within its panel for visual comparability. We overlay four signals across generations: Track II drift $\Delta U_{\mathrm{LLN,cov}}(\delta)$, Track I drift $\Delta G_{\mathrm{KF}}(\delta)$, distinct-2, and hashed $n$-gram HHI. Left: S2 (true recursion). Right: S1 (restart-from-base). The key qualitative trend is that the large early Track-II contraction in S2 precedes (and is accompanied by) stronger surface-form concentration (rising HHI) than in S1.
What the overlay reveals.

### v2

@@ Appendix A Mathematical Deductions and Proof of Theorems
$\left\lVert\hat{C}_{t}-C\right\rVert_{op}=O_{p}(t^{-1/2})$ (A.12)
@@
$\left\lVert\Delta_{t}\right\rVert_{op}\leq\left\lVert\hat{C}_{t}^{-1}\right\rVert_{op}\left\lVert\hat{C}_{t}-C\right\rVert_{op}\left\lVert C^{-1}\right\rVert_{op}.$ (A.14)
@@
This appendix connects Theorems 12 to the two computable Sigma-UB diagnostics in Section 3.4 and justifies the normalizations used for online monitoring.
@@ B.1 Regularized $\log$-determinant and PSD monotonicity
since $X+\delta I_{m},Y+\delta I_{m}\in\mathbb{S}^{m}_{++}$ for any $\delta>0$ and $\log\det(\cdot)$ is Loewner-monotone on $\mathbb{S}^{m}_{++}$. This is the basic tool that turns PSD majorants into one-sided envelopes for the (regularized) log-volume.
B.2 Bridge I: deterministic envelope via a tail-energy budget (Track I)
B.2.1 From Theorem 1 to a PSD majorant
Recall $G^{(k)}=G_{A}^{(k)}+G_{B}^{(k)}$ with $G_{B}^{(k)}\succeq 0$. By Weyl’s inequality,
$\lambda_{i}\!\left(G^{(k)}\right)\leq\lambda_{i}\!\left(G_{A}^{(k)}\right)+\lambda_{1}\!\left(G_{B}^{(k)}\right)\qquad(i=1,\dots,m).$ (B.2)
Let $\beta_{k}:=\lambda_{\max}(G_{B}^{(k)})=\lambda_{1}(G_{B}^{(k)})$. Then, in PSD order,
$G^{(k)}\preceq G_{A}^{(k)}+\beta_{k}I_{m}.$ (B.3)
Shifting by $\delta I_{m}$ and applying (B.1) gives the (incomputable) envelope
$\log\det\!\big(G^{(k)}+\delta I_{m}\big)\leq\log\det\!\big(G_{A}^{(k)}+(\beta_{k}+\delta)I_{m}\big).$ (B.4)
B.2.2 Replacing $\beta_{k}$ by a computable budget $\widehat{\beta}_{k}$
Because $G_{B}^{(k)}\succeq 0$, we have the bound
$\beta_{k}=\lambda_{\max}(G_{B}^{(k)})\leq\mathrm{tr}(G_{B}^{(k)}).$ (B.5)
@@ B.2.2 Replacing $\beta_{k}$ by a computable budget $\widehat{\beta}_{k}$
Combining (B.3) and (B.7) yields the computable PSD majorant
@@
B.2.3 Valid regularized envelope and the normalized gain form
Define the computable envelope
$\mathcal{U}_{\mathrm{KF}}^{(k)}(\delta):=\log\det\!\left(G_{A}^{(k)}+(\widehat{\beta}_{k}+\delta)I_{m}\right).$ (B.9)
Then (B.8) and (B.1) imply
$\log\det\!\big(G^{(k)}+\delta I_{m}\big)\leq\mathcal{U}_{\mathrm{KF}}^{(k)}(\delta).$ (B.10)
To interpret this envelope across checkpoints, we normalize out the isotropic baseline $(\widehat{\beta}_{k}+\delta)I_{m}$, whose log-determinant equals $m\log(\widehat{\beta}_{k}+\delta)$. Define
$\mathcal{G}_{\mathrm{KF}}^{(k)}(\delta)$ $:=\mathcal{U}_{\mathrm{KF}}^{(k)}(\delta)-m\log(\widehat{\beta}_{k}+\delta).$ (B.11)
Equivalently,
$\mathcal{G}_{\mathrm{KF}}^{(k)}(\delta)$ $=\log\det\!\left(I_{m}+\frac{1}{\widehat{\beta}_{k}+\delta}G_{A}^{(k)}\right)=\sum_{i=1}^{m}\log\!\left(1+\frac{\lambda_{i}(G_{A}^{(k)})}{\widehat{\beta}_{k}+\delta}\right).$ (B.12)
Thus, $\mathcal{G}_{\mathrm{KF}}^{(k)}(\delta)$ defines a dimensionless gain relative to the tail-budget scale: it decreases as the eigenvalues of $G_{A}^{(k)}$ contract, and it is most sensitive when eigenvalues are near the $\delta$ floor.
B.2.4 Gap control relative to the ideal (incomputable) envelope
Define the ideal but generally incomputable envelope as
$\mathcal{U}^{(k)}_{\mathrm{KF},\star}(\delta):=\log\det\!\left(G_{A}^{(k)}+(\beta_{k}+\delta)I_{m}\right).$
Since $\widehat{\beta}_{k}\geq\beta_{k}$, by monotonicity of the log-determinant,
@@ B.2.4 Gap control relative to the ideal (incomputable) envelope
This implies
@@
Thus, the normalized quantity $\mathcal{G}_{\mathrm{KF}}^{(k)}(\delta)$ remains uniformaly controlled by the (unknown) ideal normalization, even when $\widehat{\beta}_{k}$ is a loose bound.
B.3 Bridge II: LLN scaling and covariance normalization (Track II)
B.3.1 LLN scaling and cancellation of unknown constants by drift
Motivated by the law-of-large-numbers (LLN) scaling suggested by Theorem 2, define
@@ B.3.1 LLN scaling and cancellation of unknown constants by drift
When the stochastic bound is written in logarithmic form as
$\log\det(G^{(k)})\;\leq\;\log K+\mathcal{U}_{\mathrm{LLN}}^{(k)}(0)\quad\text{for an unknown finite }K,$
then baseline differencing removes $\log K$. Accordingly, define the differenced quantity
@@
B.3.2 Covariance-normalized form
Define the covariance-normalized statistic
$\mathcal{U}_{\mathrm{LLN,cov}}^{(k)}(\delta)$ $:=\mathcal{U}_{\mathrm{LLN}}^{(k)}(\delta)-m\log n_{k}=\log\det(G_{A}^{(k)}+\delta I_{m})-m\log n_{A}.$ (B.17)
Equivalently,
$\mathcal{U}_{\mathrm{LLN,cov}}^{(k)}(\delta)=\log\det\!\left(\frac{1}{n_{A}}G_{A}^{(k)}+\frac{\delta}{n_{A}}I_{m}\right).$ (B.18)
Under the i.i.d. assumption, $\frac{1}{n_{A}}G_{A}^{(k)}\to C$ as $n_{A}\to\infty$, so for large $n_{A}$,
$\mathcal{U}_{\mathrm{LLN,cov}}^{(k)}(\delta)\approx\log\det(C),$
up to the vanishing regularization $\delta/n_{A}$. Hence $\mathcal{U}_{\mathrm{LLN,cov}}^{(k)}(\delta)$ is size-invariant and is the natural object to difference over checkpoints for collapse monitoring.
@@ B.4 Numerical computation of $\log$-determinant
Appendix C Supplementary Experiments, Reproducibility, and Extended Diagnostics (TECH)
C.1 Section Roadmap
This appendix section has three goals: (i) to make the TECH recursion loop fully reproducible from first principles, (ii) to make all main-text plots and scalars auditable from exported tables, and (iii) to provide extended diagnostics that contextualize geometric contraction.
(i) Reproducibility from raw sources to Sigma-UB curves.
The recursion loop is sensitive to implementation choices. We therefore document the entire pipeline for the reported TECH run: (a) how the real TECH corpus and frozen prompt bank are constructed, (b) how synthetic pools are regenerated deterministically each generation, (c) how training mixtures are formed under a constant token budget, and (d) how evaluation responses, embeddings, and Sigma-UB metrics are computed.
(ii) Plot–table consistency and “single source of truth.”
All scalar summaries reported in the main text are computed directly from exported per-generation tables. To prevent transcription errors, we include the exact plotted points as tables in Appendix C.4.
(iii) Extended diagnostics.
Because Sigma-UB is a geometric diagnostic computed from embedding Gram structure, we also report two surface-form baselines on the same generated outputs used to form the Sigma-UB embeddings (Appendix C.5). These baselines are included for interpretability: they quantify lexical repetition directly, and help distinguish cases where geometric contraction aligns with (or departs from) visible degeneracy in text.
Baseline A: Distinct-$2$ (Li et al., 2016). Distinct-$2$ is a standard diversity score defined as the ratio
$\mathrm{Distinct}\text{-}2\;=\;\frac{\#\{\text{unique bigrams}\}}{\#\{\text{total bigrams}\}},$
computed on the generated corpus (or per-generation batch). Lower Distinct-$2$ indicates that the model is reusing the same short phrases more frequently, i.e., increased short-range repetition.
Baseline B: Hashed $n$-gram HHI (Cormode and Muthukrishnan, 2005; Itagaki et al., 2007) (global concentration of repeated phrases). To capture longer-range and heavier-tailed repetition, we compute a Herfindahl–Hirschman Index (HHI) over $n$-gram frequencies: if $c_{j}$ is the count of $n$-gram type $j$ and $p_{j}=c_{j}/\sum_{\ell}c_{\ell}$, then
$\mathrm{HHI}_{n}\;=\;\sum_{j}p_{j}^{2}.$
Higher HHI implies that a small number of $n$-grams account for a large share of the text, i.e., strong concentration and mode collapse at the surface level.
Finally, we provide an optional bucket-localization slice using the frozen TECH prompt taxonomy (Appendix C.6) to identify where contraction is most pronounced. Importantly, these surface-form baselines are contextual, not definitional: they are reported to help interpret when geometric contraction coincides with familiar repetition patterns, without redefining collapse in terms of lexical features.
C.2 TECH controlled assets: data provenance, licensing, and construction
This paper’s experiments rely on two frozen assets for TECH: (i) a real corpus $R$ of fixed-length contexts used as prompts for synthetic regeneration and as the real-data component of mixtures, and (ii) a frozen prompt bank $P$ used for evaluation.
Real TECH corpus $R$.
We collect TECH documents from permissively licensed software and technical documentation sources (e.g., The Rust Book (MIT/Apache-2.0), Kubernetes documentation (CC BY 4.0), PostgreSQL documentation (BSD-style), SQLite documentation (Public Domain), and Python documentation (PSF license)), verifying license terms per source/page. Documents are converted to plain text with boilerplate removal and code-block stripping, filtered to English, and tokenized using the OPT-125M GPT-2 BPE tokenizer to match the base model’s vocabulary. We then extract 64-token contexts using a sliding window with fixed stride and enforce minimum-length and quality filters. Finally, we remove exact duplicates (content hashes) and near-duplicates prior to selecting the final $N=1000$ contexts.
Frozen TECH prompt bank $P$ (stable prompt IDs and buckets).
Independently of $R$, we construct a TECH prompt bank designed to elicit a range of response regimes (open-ended creative prompts through neutral factual prompts) while maintaining evaluation stability across checkpoints. Prompts receive persistent prompt IDs and are categorized into buckets such as Creative, Divergent, Analogy/ELI5, What-if, and Neutral.
C.3 Reproducibility specification for S1 and S2
@@
We summarize every operational choice that affects outcomes: model hyperparameters; synthetic regeneration and mixture; decoding; seeds; Sigma-UB computation; evaluation metrics; hardware/software; and the formal difference between S1 and S2. Unless explicitly stated, S1 and S2 share the same configuration; the only intended difference is whether weights are reset to $M^{(0)}$ (S1) or carried forward $M^{(g-1)}$ (S2).
@@
Table 2: Model and training hyperparameters. These hyperparameters are held fixed across S1 and S2 so that observed differences are attributable to data recursion vs. data+weight recursion, rather than to optimization budget changes.
@@
Generations reported $g=0,\dots,50$ (51 checkpoints) Main text TECH run shown in Figures 12
@@
Table 3: Data generation and mixing. We use constant-budget mixtures and, for the reported runs, set $\alpha=0$ to isolate synthetic recursion dynamics without real-data refresh.
@@
Max tokens (evaluation) 256 Prompt-bank responses Used to obtain sufficiently long responses
for embedding/geometry
Table 4: Text generation (decoding) parameters. Decoding is frozen across checkpoints to prevent evaluation drift and to keep the generation distribution comparable over time.
@@
Table 5: Sigma-UB computation. Both tracks are computed from the same frozen embeddings and the same observed index set, so drift is attributable to model evolution rather than to evaluation sampling.
@@
Table 6: Evaluation metrics. The appendix focuses on geometry diagnostics and lightweight surface-form proxies; perplexity-based baselines are excluded by design.
@@
Table 7: Computational environment. Hardware does not change the definition of metrics but can affect throughput; determinism is enforced via fixed seeds and fixed decoding.
@@
Table 8: Prompt buckets for localization. Bucket IDs and assignments are fixed by DecayBench; slicing by bucket does not change the evaluation distribution over time, it only reveals where contraction is concentrated within a frozen taxonomy.
Operational end-to-end loop (implementation-level clarification).
For auditability, we summarize the precise loop that produces one point on each drift curve: (1) regenerate $S^{(g)}$ by prompting each $x\in R$ and sampling a fixed-length continuation with fixed decoding and per-context deterministic seeds; (2) form $D^{(g)}(\alpha)$ under a constant block budget of $|R|$; (3) fine-tune under either S1 (initialize from $M^{(0)}$) or S2 (initialize from $M^{(g-1)}$) using fixed hyperparameters; (4) evaluate on the frozen prompt bank $P$ using fixed decoding and per-prompt deterministic seeds to produce $n_{k}$ responses; (5) embed responses with the frozen encoder, select a fixed observed index set $\mathcal{I}_{A}$ once per run, and compute both Sigma-UB tracks and (optionally) surface proxies. These steps are deterministic given the experiment card above and the fixed input assets.
C.4 Exact per-generation values for the main-text drift plots
In particular, the S1 vs. S2 summary in Table 1 is computed directly from these per-generation series (including the baseline point $g=0$, where drift is 0 by definition).
@@ C.4 Exact per-generation values for the main-text drift plots
C.5 Surface-form proxy diagnostics and alignment with Sigma-UB
Sigma-UB is specifically constructed to follow the geometry of the gram matrix instead of merely capturing surface-level repetition, while still providing robust statistical validity and support for inference. Nevertheless, it is useful to ask whether geometry contraction coincides with familiar surface-form baseline metrics. We therefore compute two lightweight proxies on the same generated text used for Sigma-UB embeddings: (i) distinct-2 (unique/total bigrams; higher is more diverse), and (ii) hashed $n$-gram HHI (higher is more concentrated/repetitive). These proxies are not collapse definitions and can be noisy, but they provide an operational sanity check.
Figure 3: Normalized overlay of Sigma-UB drifts and surface-form proxies (TECH). Each curve is min–max normalized to $[0,1]$ within its panel for visual comparability. We overlay four signals across generations: Track II drift $\Delta U_{\mathrm{LLN,cov}}(\delta)$, Track I drift $\Delta G_{\mathrm{KF}}(\delta)$, distinct-2, and hashed $n$-gram HHI. Left: S2 (true recursion). Right: S1 (restart-from-base). The key qualitative trend is that the large early Track-II contraction in S2 precedes (and is accompanied by) stronger surface-form concentration (rising HHI) than in S1.